# DArFace: Deformation Aware Robustness for Low Quality Face Recognition

### v3

@@ 3.2 Preliminaries
1:  Input: Dataset $\mathcal{D}=\{(x_{n},y_{n})\}_{n=1}^{N}$, model $F_{\theta}$, transformation parameters $\psi=(\varphi,\Delta u,\Delta v,\lambda,\alpha,\sigma)$, learning rate $\eta$, adversarial step size $\eta_{\psi}$, number of ascent steps $K$, projection bounds $\mathcal{C}$, trade-off coefficient $\lambda_{\text{cont}}$
2:  Output: Model $F_{\theta}$
3:  for $j=0,\dots,T-1$ do
4:     Sample mini-batch $\{(x,y)\}\sim\mathcal{D}$
5:     Initialize transformation parameters: $\psi\leftarrow\psi_{0}$
6:     for $k=1$ to $K$ do
7:        Compute loss: $\mathcal{L}_{\text{trans}}=\mathcal{L}(F_{\theta}(T_{\psi}(x)),y)$
8:        Compute gradient: $g_{\psi}=\nabla_{\psi}\mathcal{L}_{\text{trans}}$
9:        Gradient ascent update: $\psi_{k+1}\leftarrow\psi_{k}+\eta_{\psi}\text{sgn}(g_{\psi})$
10:        Project to valid range: $\psi^{*}\leftarrow\text{Proj}_{\mathcal{C}}(\psi_{k+1})$
11:     end for
12:     Set $\psi^{*}\leftarrow\psi$
13:     Compute adversarial loss: $\mathcal{L}_{\text{trans}}=\mathcal{L}(F_{\theta}(T_{\psi^{*}}(x)),y)$
14:     Compute clean loss: $\mathcal{L}_{\text{clean}}=\mathcal{L}(F_{\theta}(x),y)$
15:     Split into clean, global, and local images:
$(X_{\text{clean}},X_{\text{global}},X_{\text{local}})=\text{split\_batch}(x)$
16:     Forward pass through backbone for all images:
$(F_{\text{clean}},F_{\text{global}},F_{\text{local}})=F_{\theta}(X_{\text{clean}},X_{\text{global}},X_{\text{local}})$
17:     Compute contrastive loss between global and local embeddings:
18:     Compute total loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{clean}}+\mathcal{L}_{\text{trans}}+\lambda_{\text{cont}}\cdot\mathcal{L}_{\text{cont}}$
19:     Update model: $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}_{\text{total}}$
20:  end for
@@ 3.3 Deformation-Aware Augmentation
Table 2: Performance comparison on mixed-quality inputs (IJB-B and IJB-C) using a ResNet-100 backbone and MS1MV2 and MS1MV3 training dataset. DArFace consistently outperforms prior baselines, particularly in low-FAR and identification metrics, highlighting its robustness against quality degradations. Note that, ArcFace+CFSM, a GAN-based approach, relies on synthetic data generation and careful tuning, making its training process more complex.
@@
[…]
@@ 4.1 Datasets
Training Dataset. We trained our models on MS1MV2 [6], MS1MV3 [8] and VGGFace2 [1] datasets. The first two have been collected and cleaned from the original MS1M (MegaFace) dataset[11]. MS1MV2 contains around 5.8M images belonging to around 85K identities, while MS1MV3 contains around 5.1M images and 93K identities. VGGFace2 contains roughly 3.1M images spanning 9K identities. Following standard facial recognition pipelines, all the images are aligned and resized to a resolution of $112\times 112$.
Evaluation Dataset. For evaluation we consider three quality categories: low, medium and high. The low quality category includes TinyFace [3] comprising of 169,403 images across 5,139 identities. The average image size is $20\times 16$ pixels. The medium quality comprises of IJB-C and IJB-B [49] datasets. IJB-B comprises of 21.8K images and 7K videos (55K frames) representing 1,845 identities. IJB-C an extension of IJB-B contains 31.3K images and 117.5K frames belonging to 3,531 identities. The high quality datasets include LFW [16], CFP-FP [38], AgeDB30 [33], CALFW [57] and CPLFW [58].
4.2 Implementation Details
We use ResNet100 [23] backbone, a standard choice in many facial recognition systems. We use SGD as the optimizer, with an initial learning rate of 0.2, momentum of 0.9 and the weight decay 0.0005, the learning rate scheduler is PolynomialLR. The model is trained for 28 epochs using the ArcFace [6] loss, however our training framework can also be used with other loss functions. The batch size per GPU is 512 and Quadro RTX 8000 are used for training.
For deformation-aware augmentation, parameters are sampled from Gaussian distributions: elasticity strength $\alpha\sim\mathcal{N}(1,0.1)$, smoothing factor $\sigma\sim\mathcal{N}(10,5)$, scale $s\sim\mathcal{N}(1,0.01)$, rotation $\theta\sim\mathcal{N}(0,0.01)$, and translation $t\sim\mathcal{N}(0,0.01)$. To ensure stability, $\alpha\in[0.05,0.25]$ and $\sigma\in[20.0,35.0]$. We set $\lambda_{\text{cont}}=0.7$. The best performance was achieved using 75% clean, 15% globally transformed, and 10% locally deformed images. Following [21], we also applied crop, low-resolution, and photometric augmentations.
4.3 High-Quality Images
@@
Table 3: Performance comparison on low-quality (TinyFace) dataset using a ResNet-100 backbone and MS1MV2, MS1MV3 and VGGFace2 training dataset. Results on TinyFace validate the effectiveness of our proposed global-local deformation strategy under extreme low-quality conditions.
@@ 4.4 Mixed-Quality Images
We evaluate the performance of our method on mixed-quality datasets which include variations in pose, illumination and resolution. To ensure the generality of the improvements, we train DArFace on two different datasets MS1MV2, MS1MV3 and compare them against the state-of-the-art methods. The performance is reported using multiple metrics, including TAR@FAR(True Acceptance Rate at various False Acceptance Rates) and Identification Accuracy at Rank 1 and Rank 5.
@@ 4.5 Low Quality Images
To assess the robustness of our method under extreme quality degradations, we evaluate our method on TinyFace dataset, which contains low quality surveillance style face images. We train our model on three different datasets MS1MV2, MS1MV3 and VGGFace2 datasets and report the identification accuracy at Rank 1 and Rank 5.
As shown in Table. 3, DArFace achieves the best performance across all the training sets. On MS1MV2 it achieves 68.24% rank 1 accuracy and 72.26% rank 5. On MS1MV3 the gains are 69.39% at rank 1 and 73.00% at rank 5. Similarly on VGGFace2 Rank-1 accuracy is 65.42% while Rank-5 is 70.70%. Compared to ArcFace and ARoFace, DArFace yields consistent gains, demonstrating that the model benefits from explicitly learning robustness to low-quality distortions during training.

### v4

@@ 3.2 Preliminaries
1:Input: Dataset $\mathcal{D}$, model $F_{\theta}$, transformation parameters $\psi_{0}$, learning rate $\eta$, adversarial step size $\eta_{\psi}$, number of ascent steps $K$, projection bounds $\mathcal{C}$, trade-off coefficient $\lambda_{\text{cont}}$
2:for mini-batch $(x,y)\sim\mathcal{D}$ do
3:  Adversarial search: $\psi\leftarrow\psi_{0}$
4:  for $k=1$ to $K$ do
5:   $\psi\leftarrow\text{Proj}_{\mathcal{C}}\!\left(\psi+\eta_{\psi}\operatorname{sgn}\big(\nabla_{\psi}\mathcal{L}(F_{\theta}(T_{\psi}(x)),y)\big)\right)$
6:  end for
7:  Losses:
8:  $\mathcal{L}_{\text{clean}}\leftarrow\mathcal{L}(F_{\theta}(x),y)$, $\mathcal{L}_{\text{trans}}\leftarrow\mathcal{L}(F_{\theta}(T_{\psi}(x)),y)$
9:  Form global/local views $X_{{g}},X_{{l}}$; get features $F_{{g}}=F_{\theta}(X_{{g}})$, $F_{{l}}=F_{\theta}(X_{{l}})$
10:  $\mathcal{L}_{\text{cont}}\leftarrow 1-\text{cos\_sim}(F_{{g}},F_{{l}}).\text{mean()}$
11:  Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\!\big(\mathcal{L}_{\text{clean}}+\mathcal{L}_{\text{trans}}+\lambda_{\text{cont}}\mathcal{L}_{\text{cont}}\big)$
12:end for
@@ 3.3 Deformation-Aware Augmentation
ArcFace [6] WebFace4M 42.19 87.60 94.43 95.12 96.83 69.03 89.40 95.70 96.03 97.50
ARoFace [36] WebFace4M 44.88 91.82 95.43 95.78 97.41 87.90 94.58 96.85 96.99 98.06
AdaFace [21] WebFace4M 44.48 92.26 96.03 96.26 97.68 90.43 95.34 97.39 97.52 98.35
DArFace (Ours) WebFace4M 44.33 92.06 95.45 95.79 97.43 88.93 94.76 96.82 97.10 98.05
ArcFace [6] WebFace12M 41.96 88.18 95.24 95.45 97.15 74.25 91.02 96.46 96.45 97.70
ARoFace [36] WebFace12M 44.61 93.05 96.14 96.09 97.48 89.51 95.62 97.35 97.25 98.15
AdaFace [21] WebFace12M 47.49 93.13 96.30 96.28 97.72 89.47 95.94 97.54 97.56 98.38
DArFace (Ours) WebFace12M 48.37 92.73 96.23 96.25 97.59 86.34 95.23 97.41 97.32 98.19
Table 2: Performance comparison on mixed-quality inputs (IJB-B and IJB-C) using a ResNet-100 backbone and MS1MV2, MS1MV3, WebFace4M and WebFace12M training dataset. DArFace consistently outperforms prior baselines, particularly in low-FAR and identification metrics, highlighting its robustness against quality degradations. Note that, ArcFace+CFSM, a GAN-based approach, relies on synthetic data generation and careful tuning, making its training process more complex.
@@
where $\text{cos\_sim}(\cdot,\cdot)$ denotes the cosine similarity between two feature vectors. A high similarity (i.e. low contrastive loss) indicates that the network preserves identity information across different deformations. The total loss used to train the network is a weighted combination of loss on clean samples, the adversarial loss on deformed samples, and the contrastive loss:
@@ 4.1 Datasets
Training Dataset. We trained our models on MS1MV2 [6], MS1MV3 [8], VGGFace2 [1], WebFace4M and WebFace12M [59] datasets. MS1MV2 contains around 5.8M images belonging to around 85K identities, while MS1MV3 contains around 5.1M images and 93K identities. VGGFace2 contains roughly 3.1M images spanning 9K identities. WebFace12M contains 12.7M images and 617K identities. WebFace4M is a subset of WebFace12M with 4.2M images belonging to 205K identities.
Evaluation Dataset. For evaluation we consider three quality categories: low, mixed and high. The low quality category includes TinyFace dataset [3], the mixed quality comprises of IJB-C and IJB-B [49] datasets. The high quality datasets include LFW [16], CFP-FP [38], AgeDB30 [33], CALFW [57] and CPLFW [58].
@@
ArcFace [6] WebFace4M 69.68 73.28
ARoFace [36] WebFace4M 71.62 74.67
AdaFace [21] WebFace4M 72.02 74.52
DArFace (Ours) WebFace4M 72.61 75.59
ArcFace [6] WebFace12M 70.25 73.20
ARoFace [36] WebFace12M 72.71 75.61
AdaFace [21] WebFace12M 72.29 74.97
DArFace (Ours) WebFace12M 73.25 76.01
Table 3: Performance comparison on low-quality (TinyFace) dataset using a ResNet-100 and MS1MV2, MS1MV3 VGGFace2, WebFace4M and WebFace12M training datasets. Results on TinyFace validate the effectiveness of our proposed global-local deformations under extreme low-quality conditions.
4.2 Implementation Details
We employ a ResNet-100 backbone with ArcFace loss [6], optimized using SGD (initial LR = 0.2, weight decay = 5e45e–4) and a PolynomialLR scheduler for 28 epochs. The framework is compatible with other loss functions.
For deformation-aware augmentation, parameters are sampled from Gaussian distributions: $\alpha\sim\mathcal{N}(1,0.1)$, $\sigma\sim\mathcal{N}(10,5)$, $s\sim\mathcal{N}(1,0.01)$, $\theta\sim\mathcal{N}(0,0.01)$, and $t\sim\mathcal{N}(0,0.01)$, with stability bounds $\alpha\in[0.05,0.25]$ and $\sigma\in[20.0,35.0]$. We set $\lambda_{\text{cont}}=0.7$ and use 75% clean, 15% globally transformed, and 10% locally deformed samples. Following [21], we also applied crop, low-resolution, and photometric augmentations.
4.3 High-Quality Images
@@ 4.4 Mixed-Quality Images
We evaluate the performance of our method on mixed-quality datasets which include variations in pose, illumination and resolution. To ensure the generality of the improvements, we train DArFace on four different datasets MS1MV2, MS1MV3, WebFace4M, WebFace12M and compare them against the state-of-the-art methods. The performance is reported using TAR@FAR(True Acceptance Rate at various False Acceptance Rates) and Identification accuracy at Rank 1 and Rank 5.
@@
Training on more challenging WebFace4M and WebFace12M datasets further demonstrate that DArFace maintains comparable or superior performance across both verification and identification metrics, reinforcing its generalization ability across varying data scales and quality levels.
@@ 4.5 Low Quality Images
To assess the robustness of our method under extreme quality degradations, we evaluate our method on TinyFace dataset, which contains low quality surveillance style face images. We train our model on five different datasets MS1MV2, MS1MV3, VGGFace2, WebFace4M and WebFace12M datasets and report the identification accuracy at Rank 1 and Rank 5.
As shown in Table. 3, DArFace achieves the best performance across all the training sets. On MS1MV2 it achieves 68.24% rank 1 accuracy and 72.26% rank 5. On MS1MV3 the gains are 69.39% at rank 1 and 73.12% at rank 5. Similarly on VGGFace2 Rank-1 accuracy is 65.42% while Rank-5 is 70.70%. Compared to ArcFace and ARoFace, DArFace yields consistent gains, demonstrating that the model benefits from explicitly learning robustness to low-quality distortions during training.
When scaled to larger datasets, DArFace demonstrate consistent gains. On WebFace4M, it reaches 72.61% at rank 1 and 75.59% at rank 5, while on the more challenging WebFace12M, it achieves 73.25% and 76.01%, surpassing all prior baselines. These results confirm that DArFace effectively generalizes to large-scale data and benefits from its global-local deformation learning, which explicitly enhances robustness to severe low-quality distortions.